Translate this sentence into an equation.
72
decreased by Vanessa's age is
18
.
Use the variable
v
to represent Vanessa's age.

Answer by Theo(13342)

72 - v = 18

solve for v to get v = 72 - 18 = 54.

vanessa's age is 54.

replace v with 54 to get 72 - 54 = 18

this confirms the solution is correct.

the solution is that the equation is 72 - v = 18.